(52+6*sq root43)power3/2+ (52-6*sq root43)power3/2  is divided by 100,the remainder is.          

Grade:12

1 Answers

Sukhendra Reddy Rompally B.Tech Mining Machinery Engg, ISM Dhanbad
12 years ago

Dear Rahul,

consider (52+6*square root43)power1/2 = a , and (52+6*square root43)power1/2 =b...so, it is acube + bcube,...which is (a+b)(asquare + bsquare -ab),...substitute the values of a and b, and u get (a+b) ( 52 + 6*square root43 + 52 - 6*square root43 - square root of 52 square - 36*43 ) = (a+b)(104-34) = 70 ( a+b)
when u divide by 10, u get 7(a+b),..re-substitute the values of a and b,..and thats the answer

THANKS and ALL THE BEST

PLZZZ APPROVE MY ANSWER IF U LIKE IT :)